# {strip}[Uncaptioned image]

### v1

@@ 4.1. Prototype 1: “Baremetal IO”
Target app: A donut spinning on display (Figure 1(b,c)). OS capabilities: Prototype 1 is a baremetal appliance for a single application. Its architecture is similar to that of microcontroller like ESP32. It demonstrates the concept of software abstracting hardware details and introduces the idea of virtualization through the timer driver.
Framebuffer a first-class IO
Unlike many education OSes where display was introduced much later if at all, we support the framebuffer as a first-class IO device. This is done via the Pi3’s hardware mailbox interfaces. Output-only, polling-only UART for debug For debugging messages, the prototype implements a minimalist UART driver. For simplicity, UART writes are synchronous and poll for completion. We found interrupt-driven UART writes unnecessarily complicated the kernel. For instance, such writes would need a ring buffer for outstanding writes, which itself needs locks; on the other hand, the lock code may print debugging messages via UART, creating a circular dependency. Thus, we keep UART writes always synchronous throughout prototypes 1–5. We defer interrupt-driven (asynchronous) inputs to USB keyboards (Prototype 4) and asynchronous output to sound (DMA, Prototype 5), for which asynchrony is absolutely necessary. Timers for periodic rendering. Prototype 1 drives per-core ARM generic timers as well as the SoC-level timer. Atop one physical timer, Prototype 1 creates multiple virtual timers.
@@ 4.2. Prototype 2: “Multitasking”
Target app: Multiple concurrent donuts, each spinning at its own pace. New OS capabilities: It introduces multitasking, runs predefined tasks, one for a donut. It reinforces the concept of virtualization, with the case of CPU.
See tasks run as donuts spin
@@
Much of the complexity comes from context switch: manipulation of CPU registers around the interrupt handlers and scheduler. The scheduler itself is simple (a single runqueue), sufficient to manage less than several tasks on a single core. All code runs at EL1 with no kernel/user separation. Everything operates on physical addresses, with memory for tasks statically reserved. The ”kernel” is a monolithic program, and apps are simply functions compiled into the same binary.
@@
Target app: “Mario” without inputs (see Section 3 for description). New OS capabilities: this prototype introduces user/kernel separation, which allows to compile/link/execute Mario independent of the kernel.
Virtual Memory
@@
VM is not enough for separation
To motivate a userspace more decoupled of the kernel, this prototype supports “infant” apps that run in their own address space and at EL0, but are still compiled and linked as part of the kernel. These infant apps are launched via jumps from the kernel code, rather than exec(). This is achieved by ensuring kuser code only issues PC-relative memory references. Evidently, such infant apps are brittle because absolute addresses would refer to the kernel address space and cause exceptions. They are only suitable for simple ”hello world” code consisting of tens of lines. This limitation motivates linking and loading Mario (around 2K lines of code) independently of the kernel.
Syscalls without files
One important decision is omission of files, deferring them to the next prototype, as file support would introduce a myriad of data structures (e.g. locks and inodes) and code, complicating the task management and syscall paths. Without files, our syscalls are just several: fork() and exit() for task lifecycles, sbrk() for Mario’s pixel buffer allocation, sleep() for timed rendering, and write() hardwired to UART for debugging. A special file-less exec() is needed due to lack of files. Our build scripts bundles Mario’s ELF executable as a opaque binary, with the kernel image at the link time. At run time, exec() parses the elf region in memory and loads the code/data segments into the user virtual address space. It also hardcodes the arguments (e.g., framebuffer address and geometry) as expected by Mario.
mmap for Mario’s direct rendering
Without device files, framebuffer access for rendering motivates our implementation of mmap. At the end of exec(), the kernel appends a 4KB granular mapping to Mario’s page table, covering the entire framebuffer (typically several MBs). For debugging ease, the virtual address is identity-mapped to the framebuffer’s physical address, enabling direct framebuffer access akin to Linux’s Direct Rendering Infrastructure (DRI) (paul2000introduction, ).
Leave Mario input-less
@@
Interestingly, Mario, even without inputs, continuously renders animated frames – such as flashing a coin on the title screen – and eventually transitioning into autoplay. This limitation encourages students to engage with the next prototype to enable full control of Mario.
Experience: see CPU cache in action
Framebuffer rendering provides a window into the impact of CPU cache behavior. First, the framebuffer must be mapped as cached to avoid significant FPS drop. Second, interesting yet subtle issue arises: the CPU cache must be flushed for the framebuffer region on every frame; otherwise, stale framebuffer pixels from previous frames will cause non-deterministic visual artifacts that will gradually disappear as cache lines hit the memory.
Since the user level (EL0) cannot flush the cache, a kernel request (e.g., via a syscall) is required. Ultimately, cache maintenance is crucial for reliable game performance.
@@ 4.4. Prototype 4: “Files”
Target apps: Additional NES games, slider, MusicPlayer. New OS capabilities: This prototype introduces the file abstraction, extending the user/kernel interfaces. Underneath files, it introduces a small filesysetm on ramdisk as well as new IO drivers. The asynchrony introduced by the new IOs necessitates additional concurrency mechanisms inside both the kernel and userspace.
Disk files for games, photos, and music
As such, the NES game engine can load additional ROMs as files (as opposed to the Mario ROM hardcoded in a C header for the engine). The slider app loads photo files as slides, and MusicPlayer loads music files and their album cover photos. We ported the xv6’s ext2-like filesystem (“xv6fs” for short) for its simplicity. The filesystem runs on a ramdisk; since all block reads/writes are synchronous, file read/write paths always execute in syscall contexts, easing understanding and debugging. Support for physical block devices, which must cope with storage hardware details, is deferred to Prototype 5. This decision decouples the complexity of a storage stack.
Device files for app IO
They encapsulate IO devices, including the framebuffer (/dev/fb), audio output (/dev/sb), and USB keyboard (/dev/events), as needed by the target apps. Detailed descriptions are provided below.
@@
Asynchronous IO paths for audio playback
The MusicPlayer app continuously writes sound samples to the device file (/dev/sb). The device driver then copies these samples and transfers them to the audio hardware via DMA for pulse width modulation output. Subsequently, DMA interrupts notify the driver to supply more samples. This setup forms a producer-consumer pipeline involving the app, the device driver, and the audio hardware, passing samples via ring buffers and synchronized by condition variables and locks—a classic OS design pattern. Any glitches in this pipeline will cause audio stutters, prompting debugging and fixes.
USB for Mario inputs
@@
We ported Circle/USPi (githubGitHubRsta2uspi, ), a baremetal USB stack for Pi3. The stack only requires a few kernel APIs for virtual timers, which were introduced in Prototype 1. Although USPi is the lightest USB stack we could find, it still has around 10K SLoC, with multiple layers of abstraction. We do not expect the students to fully understand and debug it. In return, the stack makes VOS extensible to more USB classes, such as ethernet adapters and mass storage, in the future.
IPC for Mario’s event loop
The app’s need for handling more than one asynchronous inputs motivate IPC. Notably, the main loop of Mario has to handle periodic timer events (for rendering frames at certain rate) and keyboard inputs. Lacking asynchronous IO and threading (available in Prototype 5), Mario shows a use case of IPC: the main loop forks two additional processes, one calling msleep() periodically and the other reading from /dev/event in a blocking manner; the two processes write events to a shared pipe (two writers), from which the main loop reads.
@@ 4.5. Prototype 5: “Desktop”
The xv6’s filesystem (“xv6fs”) becomes the bottleneck for new apps. (1) Capacity: xv6fs only supports files up to 270KB, insufficient for larger photos, videos, and DOOM’s game assets. (2) Performance: xv6fs reads and writes single blocks instead of block ranges, making it unacceptably slow for loading files in MBs. (3) Interoperability: While it is possible to mount xv6fs on SD cards, commodity OSes cannot access these partitions without custom programs to parse and manipulate xv6fs.
@@
These limitations motivate for porting a commodity filesystem to Prototype 5. We chose ChaN’s FatFS, a FAT32 implementation used by popular embedded systems (chan_fatfs, ). In around 6K SLoC, FatFS exposes high-level APIs with semantics close to our file syscalls (e.g. open, close, read/write). FatFS does not have an inode concept, encapsulating on-disk data structures behind its APIs, whereas each file or directory in xv6fs has an in-memory inode. To bridge the gap, Prototype 5 maintains pseudo-inodes for each opened FAT file and directory. Prototype 5 further interposes file syscalls and dispatches them to either xv6fs or FatFS based on paths. At runtime, the OS mounts on its root filesystem (in xv6fs) under ‘/’ and mounts the FAT32 partition under ‘/d’.
@@
To support FatFS, Prototype 5 includes a baremetal SD card driver. Unlike a full MMC (multimedia card) subsystem which typically has tens of KSLoC (mmc, ), our driver is only 600 SLoC. It initializes the controller and card, and performs synchronous reads and writes of single blocks or block ranges, polling for their completion. Although the hardware is capable of DMA, our driver does not use it.
Non-blocking IO for key-polling games
@@
, such as DOOM. This necessitates minor enhancements to the kernel IO path, including the VFS, device files, and the USB driver, to handle the non-blocking flag and peek the ring buffer without waiting for new data.
Threading for SDL audio
@@
SDL uses a dedicated thread to stream audio samples to the device file, running in parallel with the thread decoding the audio. Furthermore, structuring Mario with multiple threads is more natural than using multiple processes that communicate via pipes (as in Prototype 4). To this end, Prototype 5 implements a simple clone() with flag CLONE_VM, which shares mm structs across tasks within the kernel. To synchronize threads, Prototype 5 adds semaphore syscalls and implements user-level mutexes and condition variables atop semaphores. It also implements user-level spinlocks.
Multicore for running many game instances
@@
Running four or more Mario instances quickly saturates a single core, motivating the OS to scale to multicore. To do so, we made the following kernel modifications : (1) The kernel’s boot code releases additional cores held in the “parked” status by the Pi3 firmware; (2) It configures MMUs for multicore cache coherency; (3) It augments the scheduler runqueue and vector tables so that each core has a copy. (4) The interrupts from ARM generic timers (arm_generic_timer, ), which drive scheduler ticks, are fed to each core, whereas interrupts from all other IO are routed to core 0 for simplicity.
Window manager for desktop experiences
@@
Our window manager (WM, in around 800 SLoC) executes as a kernel thread. Compared to running WM as a user process (e.g., as in Android), our design favors simplicity by avoiding the need to implement shared memory IPC and protocols for exchanging drawing contents across process boundaries.
@@
To draw via the WM, apps render indirectly to a device file (/dev/surface). Internally, WM maintains a list of such surfaces drawn by concurrent apps and periodically composites these surfaces onto the hardware framebuffer. Additionally, the WM tracks the “focused” app window and dispatches input events only to this app, which are returned to the app’s read() of /dev/event1. To display many apps simultaneously, WM can render overlapped windows and tracks their z-order. For efficiency, WM tracks dirty window regions across composition rounds and only redraws these regions. It intercepts special key combinations (e.g., ctrl+tab) for switching window focus or moving windows around. It supports floating, semi-transparent windows, allowing sysmon to stay on top of all other running apps (see Figure 1(m)).
@@ 5.1. Debugging support
Most instructional OSes rely on software emulators (e.g., QEMU) for debugging. However, this approach is inadequate to us, as certain nuanced and challenging bugs only manifest on actual hardware. (1) CPU cache behavior: Real hardware may need manual cache flushes (e.g., for framebuffer updates) due to possible incoherence without proper MMU setup. (2) Uninitialized memory: Unlike QEMU’s zeroed memory, real hardware has arbitrary values in uninitialized memory. (3) IO differences, e.g. GPU framebuffers may be mapped to arbitrary addresses on real hardware. (4) Parallelism: QEMU often lacks the non-deterministic thread interleaving seen on real hardware.
@@
  • Panic button. We reserve ARMv8’s FIQ (fast interrupt) for triggering emergency dumps. A GPIO-triggered FIQ (e.g. connected to a physical push button) stays unmasked and is routed round-robin to all cores. In case the kernel is deadlocked and IRQs are masked, button push events dump call stacks and registers from all cores.
@@
Our kernel supports low-cost JTAG debuggers (e.g., FTDI-based, Segger J-Link) via the Pi3’s GPIO pins. JTAG enables bare-metal debugging from a development machine, typically using GDB over a proxy like OpenOCD (openOCD, ) (supporting breakpoints, single-stepping, etc). However, our experience shows that JTAG, on Pi3, requires tedious setup and lacks key features, such as halting on reset, loading programs, and reading system registers. For our purposes, JTAG is less effective than self-hosted debugging described above.

### v2

@@ 4.1. Prototype 1: “Baremetal IO”
Target app: A donut spinning on display (Figure LABEL:fig:collage(b,c)). OS capabilities: Prototype 1 is a baremetal appliance for a single application. Its architecture is similar to that of a microcontroller like ESP32. It demonstrates the concept of software abstracting hardware details and introduces the idea of virtualization through the timer driver.
Unlike many education OSes where display was introduced much later if at all, we support the framebuffer as a first-class IO device via the Pi3’s hardware mailbox interfaces. For debugging messages, the prototype implements a minimalist output-only, polling-based UART driver. For simplicity, UART writes are synchronous. We found interrupt-driven UART writes unnecessarily complicated the kernel; for instance, such writes would need a ring buffer for outstanding writes, which itself needs locks; on the other hand, the lock code may print debugging messages via UART, creating a circular dependency. Thus, we keep UART writes always synchronous throughout prototypes 1–5, and defer interrupt-driven (asynchronous) inputs and outputs to USB keyboards (Prototype 4) and DMA sound (Prototype 5), respectively. Finally, Prototype 1 drives per-core ARM generic timers and a SoC-level timer, as well as multiple virtual timers above these.
@@ 4.2. Prototype 2: “Multitasking”
Target app: Multiple concurrent donuts spinning at their own pace. New OS capabilities: This prototype introduces multitasking, runs predefined tasks, one for a donut. It reinforces the concept of virtualization, with the case of CPU.
@@
Much of the complexity of enabling the target app comes from context switch: manipulation of CPU registers around the interrupt handlers and scheduler. The scheduler itself is simple (a single runqueue), sufficient to manage several tasks on a single core. All code runs at EL1 with no kernel/user separation. Everything operates on physical addresses, with memory for tasks statically reserved. The “kernel” is a monolithic program, and apps are simply functions compiled into the same binary.
@@
Target app: mario without inputs (see Section 3 for description). New OS capabilities: This prototype introduces user/kernel separation, which allows us to compile, link, and execute mario independent of the kernel.
@@
The VM is not enough for motivating the user-kernel separation. To motivate a userspace more decoupled from the kernel, this prototype supports “infant” apps that run in their own address space and at EL0, but are still compiled and linked as part of the kernel. These infant apps are launched via jumps from the kernel code, rather than exec(). This is achieved by ensuring the app code only issues PC-relative memory references. Evidently, such infant apps are brittle because absolute addresses would refer to the kernel address space and cause exceptions. They are only suitable for simple “hello world” code consisting of tens of lines. This limitation motivates linking and loading mario (around 2K lines of code) independently of the kernel.
One important decision is the omission of files, deferring them to the next prototype, as file support would introduce a myriad of data structures (e.g. locks and inodes) and code, complicating the task management and syscall paths. Without files, our syscalls are just several: fork() and exit() for task lifecycles, sbrk() for mario’s pixel buffer allocation, sleep() for timed rendering, and write() hardwired to UART for debugging. While some have questioned the relevance of fork() in production systems (10.1145/3317550.3321435, ), we found it valuable for exercising the concepts of VM in an educational context. A special file-less exec() is needed due to lack of files. Our build scripts bundle mario’s ELF executable as an opaque binary, with the kernel image at the link time. At run time, exec() parses the ELF region in memory and loads the code/data segments into the user virtual address space. It also hardcodes the arguments (e.g., framebuffer address and geometry) as expected by mario.
At the end of exec(), the kernel appends a 4KB granular mapping to mario’s page table, covering the entire framebuffer (typically several MBs). For debugging ease, the virtual address is identity-mapped to the framebuffer’s physical address, enabling direct framebuffer access akin to Linux’s Direct Rendering Infrastructure (DRI) (paul2000introduction, ).
@@
Interestingly, mario, even without inputs, continuously renders animated frames—such as flashing a coin on the title screen—and eventually transitioning into autoplay. This limitation encourages students to engage with the next prototype to enable full control of mario.
Framebuffer rendering provides a window into the impact of CPU cache behavior. First, the framebuffer must be mapped as cached to avoid significant FPS drop. Second, an interesting yet subtle issue arises: the CPU cache must be flushed for the framebuffer region on every frame; otherwise, stale framebuffer pixels from previous frames will cause non-deterministic visual artifacts that will gradually disappear as cache lines hit the memory.
Since the user level (EL0) cannot flush the cache under typical hardware configurations, a kernel request (e.g., via a syscall) is required. Ultimately, cache maintenance is crucial for reliable game performance.
Students work at both the user and kernel level to build Prototype 3. On the kernel side, they must implement key functions such as creating kernel page tables, setting up the kernel idmap, as well as completing the syscall path for exec(). On the user side, they are tasked with implementing argument parsing in the provided NES emulator and finishing the graphics buffer flushing function.
@@ 4.4. Prototype 4: “Files”
Target apps: Additional NES games, slider, MusicPlayer. New OS capabilities: This prototype introduces the file abstraction, extending the user/kernel interfaces. Underneath files, it introduces a small filesystem on ramdisk as well as new IO drivers. The asynchrony introduced by the new IOs necessitates additional concurrency mechanisms inside both the kernel and userspace.
We store games, photos and music as disk files. As such, the NES game engine can load additional ROMs as files (as opposed to the mario ROM hardcoded in a C header for the engine). The slider app loads photo files as slides, and MusicPlayer loads music files and their album cover photos. We ported xv6’s ext2-like filesystem (“xv6fs” for short) for its simplicity. The filesystem runs on a ramdisk; since all block reads/writes are synchronous, file read/write paths always execute in syscall contexts, easing understanding and debugging. Support for physical block devices, which must cope with storage hardware details, is deferred to Prototype 5. This decision decouples the complexity of a storage stack.
App IO is supported through device files. They encapsulate IO devices, including the framebuffer (/dev/fb), audio output (/dev/sb), and USB keyboard (/dev/events), as needed by the target apps. Detailed descriptions are provided below.
@@
For audio playback, we implement asynchronous IO paths. The MusicPlayer app continuously writes sound samples to the device file (/dev/sb). The device driver then copies these samples and transfers them to the audio hardware via DMA for pulse width modulation output. Subsequently, DMA interrupts notify the driver to supply more samples. This setup forms a producer-consumer pipeline involving the app, the device driver, and the audio hardware, passing samples via ring buffers and synchronized by condition variables and locks—a classic OS design pattern. If there are problems in this pipeline—such as buffer corruption or deadlocks—audio playback will stutter or stop, providing immediate and clear feedback for debugging.
@@
We ported Circle/USPi (githubGitHubRsta2uspi, ), a baremetal USB stack for Pi3. The stack only requires a few kernel APIs for virtual timers, which were introduced in Prototype 1. Although USPi is the lightest USB stack we could find, it still has around 10K SLoC, with multiple layers of abstraction. We do not expect the students to fully understand and debug it. In return, the stack makes Proto extensible to more USB classes, such as ethernet adapters and mass storage, in the future.
The app’s need for handling multiple asynchronous inputs motivates adopting IPC for the app’s event loop . Notably, the main loop of mario has to handle periodic timer events (for rendering frames at certain rate) and keyboard inputs. Lacking asynchronous IO and threading (available in Prototype 5), mario shows a use case of IPC: the main loop forks two additional processes, one calling msleep() periodically and the other reading from /dev/event in a blocking manner; the two processes write events to a shared pipe (two writers), from which the main loop reads.
In this prototype, students work on both the kernel and user sides. On the kernel side, they implement functions that enable read and write operations on device files and complete the keyboard driver. On the user side, they develop code for terminal programs such as shell, ls, and cat, and write code to invoke system calls for reading and writing device files, thereby utilizing the framebuffer and keyboard.
Figure 6. The architecture of Prototype 5
@@ 4.5. Prototype 5: “Desktop”
xv6’s filesystem (“xv6fs”) becomes the bottleneck for new apps. (1) Capacity: xv6fs only supports files up to 270KB, insufficient for larger photos, videos, and DOOM’s game assets. (2) Performance: xv6fs reads and writes single blocks instead of block ranges, making it unacceptably slow for loading files in MBs. (3) Interoperability: While it is possible to mount xv6fs on SD cards, commodity OSes cannot access these partitions without custom programs to parse and manipulate xv6fs.
@@
These limitations motivate porting a commodity filesystem to Prototype 5. We chose ChaN’s FatFS, a FAT32 implementation used by popular embedded systems (chan_fatfs, ). In around 6K SLoC, FatFS exposes high-level APIs with semantics close to our file syscalls (e.g. open, close, read/write). FatFS does not have an inode concept, encapsulating on-disk data structures behind its APIs, whereas each file or directory in xv6fs has an in-memory inode. To bridge the gap, Prototype 5 maintains pseudo-inodes for each opened FAT file and directory. Prototype 5 further interposes file syscalls and dispatches them to either xv6fs or FatFS based on paths. At runtime, the OS mounts on its root filesystem (in xv6fs) under ‘/’ and mounts the FAT32 partition under ‘/d’.
@@
To support FatFS, Prototype 5 includes a baremetal SD card driver. Unlike a full MMC (multimedia card) subsystem, which typically has tens of KSLoC (mmc, ), our driver is only 600 SLoC. It initializes the controller and card, and performs synchronous reads and writes of single blocks or block ranges, polling for their completion. Although the hardware is capable of DMA, our driver does not use it.
@@
We adopt Non-blocking IO for key-polling games, such as DOOM. This necessitates minor enhancements to the kernel IO path, including the VFS, device files, and the USB driver, to handle the non-blocking flag and peek the ring buffer without waiting for new data.
@@
We use threading to process SDL audio concurrently. SDL uses a dedicated thread to stream audio samples to the device file, running in parallel with the thread decoding the audio. Furthermore, structuring mario with multiple threads is more natural than using multiple processes that communicate via pipes (as in Prototype 4). To this end, Prototype 5 implements a simple clone() with flag CLONE_VM (a Linux-like interface chosen to ease the porting of existing user libraries); within the kernel, threads are implemented by sharing mm structs across tasks. To synchronize threads, Prototype 5 adds semaphore syscalls and implements user-level mutexes and condition variables atop semaphores. It also implements user-level spinlocks.
@@
Running four or more mario instances quickly saturates a single core, motivating the OS to scale to multicore. To do so, we made the following kernel modifications. (1) The kernel’s boot code releases additional cores held in the “parked” status by the Pi3 firmware. (2) It configures MMUs for multicore cache coherency. (3) It augments the scheduler runqueue and vector tables so that each core has a copy. (4) The interrupts from ARM generic timers (arm_generic_timer, ), which drive scheduler ticks, are fed to each core, whereas interrupts from all other IO are routed to core 0 for simplicity.
@@
We implement window manager (WM, in around 800 SLoC) to provide desktop experiences. Our window manager executes as a kernel thread. Compared to running WM as a user process (e.g., as in Android), our design favors simplicity by avoiding the need to implement shared memory IPC and protocols for exchanging drawing contents across process boundaries.
@@
To draw via the WM, apps render indirectly to a device file (/dev/surface). Internally, WM maintains a list of such surfaces drawn by concurrent apps and periodically composites these surfaces onto the hardware framebuffer. Additionally, the WM tracks the “focused” app window and dispatches input events only to this app, which are returned to the app’s read() of /dev/event1. To display many apps simultaneously, WM can render overlapped windows and tracks their z-order. For efficiency, WM tracks dirty window regions across composition rounds and only redraws these regions. It intercepts special key combinations (e.g., ctrl+tab) for switching window focus or moving windows around. It supports floating, semi-transparent windows, allowing sysmon to stay on top of all other running apps (see Figure LABEL:fig:collage(m)).
In this prototype, students again work on both kernel and user code. On the kernel side, they implement system calls needed for the design, such as completing clone() for threading and modifying open() to support non-blocking I/O. They also finish the SD card driver and FAT filesystem to enable file system support. On the user side, students complete applications that leverage these new kernel features. For example, in the MusicPlayer app, they add a clone() call to create a worker thread that streams audio samples to the device file.
@@ 5.1. Debugging support
Most instructional OSes rely on software emulators (e.g., QEMU) for debugging. However, this approach is inadequate to us, as certain nuanced and challenging bugs only manifest on actual hardware. (1) CPU cache behavior. Real hardware may need manual cache flushes (e.g., for framebuffer updates) due to possible incoherence without proper MMU setup. (2) Uninitialized memory. Unlike QEMU’s zeroed memory, real hardware has arbitrary values in uninitialized memory. (3) IO differences. E.g. GPU framebuffers may be mapped to arbitrary addresses on real hardware. (4) Parallelism. QEMU often lacks the non-deterministic thread interleaving seen on real hardware. To address these issues, Proto supports debugging on both QEMU and real hardware.
@@
  • Panic button. We reserve ARMv8’s FIQ (fast interrupt) for triggering emergency dumps. A GPIO-triggered FIQ (e.g., connected to a physical push button) remains unmasked at all times. When the button is pressed, the FIQ is routed in a round-robin fashion so that each occurrence is handled by a different core. This mechanism ensures that if the kernel is deadlocked and IRQs are masked, pressing the button will dump call stacks from all cores.
@@
Our kernel supports low-cost JTAG debuggers (e.g., FTDI-based, Segger J-Link) via the Pi3’s GPIO pins. JTAG enables baremetal debugging from a development machine, typically using GDB over a proxy like OpenOCD (openOCD, ) (supporting breakpoints, single-stepping, etc). However, our experience shows that JTAG on Pi3 requires tedious setup and lacks key features, such as halting on reset, loading programs, and reading system registers. For our purposes, JTAG is less effective than self-hosted debugging described above.